Speed of wave travelling along a stretched string may depend on
a)tension in the string
b)length of the string
Derive the formula for speed of the wave along a stretched string using dimensions?
1
Expert's answer
2019-10-21T09:17:18-0400
"v=T^a\mu^b"

"[v]=[T]^a[\mu]^b"

"LT^{-1}=(MLT^{-2})^a(ML^{-1})^b"

"-1=-2a\to a=\frac{1}{2}"

"a=-b\to b=-\frac{1}{2}"

Thus,


"v=T^{\frac{1}{2}}\mu^{-\frac{1}{2}}=\sqrt{\frac{T}{\mu}}"